OpenStudy (anonymous):

HELP! Medal & Fan :) The sum of the angles in a convex 10-sided polygon is: 1800° 188° 3600° 1440°

OpenStudy (campbell_st):

the formula for angle sum of a polygon is \[Angle~Sum = (n - 2) \times 180\] where n = number of sides

OpenStudy (campbell_st):

just substitute n = 10 and calculate the answer

OpenStudy (anonymous):

And you will get 1440 i think